An object is thrown upward at a speed of 91 feet per second by a machine from a height of 15 feet off the ground. The height h of the object after t seconds can be found using the equation h = − 16t^2 + 91t + 15
(a) When will the height be 102 feet?
(b) When will the object reach the ground?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

An object is thrown upward at a speed of 91 feet per second by a machine from a height of 15 feet off the ground. The height h of the object after t seconds can be found using the equation h = − 16t^2 + 91t + 15
When will the height be 102 feet?
h = − 16t^2 + 91t + 15 = 102
Solve for t
=======================
When will the object reach the ground?
h = − 16t^2 + 91t + 15 = 0
Solve for t

Answer by ikleyn(52778)

.

(a)  It will happen, when h(t) = 102,  i.e.


     -16t^2 + 91t + 15 = 102.


     It is your basic equation to start.

     Solve it by any way you want.


     You will get 2 (two) roots.


     One root is the solution when the object moves up;  the other root is the solution, when the object falls down.



(b)  It will reach the ground, when h(t) = 0, i.e.


     -16t^2 + 91t + 15 = 0.


     So, this time you should solve this equation.

Happy solving (!)